OpenStudy (anonymous):

John was competing in a kayaking competition and he traveled 7 miles upstream and 10 miles downstream with a current of 2 miles per hour. With an average of "x" miles per hour, what is the amount of time elapsed in one round trip?

OpenStudy (phi):

rate * time = distance they tell you the rate is x miles per hour the distance of a round trip is 7 up and 10 down= 17 miles total x * time = 7+10 x * time = 17 solve for time

OpenStudy (anonymous):

its x to the power of time = 17?

OpenStudy (phi):

if you had 2 * 3 = 6 but you did not know it was 3, so you had 2 * a = 6 do you see if you divide both sides by 2, you can find a ? do the same to find time in x * time = 17